OpenStudy (anonymous):

Find the sum negative two-eighteenths plus two-thirds. Be sure your answer is reduced to simplest form. negative two-ninths negative five-eighths zero positive five-ninths

OpenStudy (ranga):

\[-\frac{ 2 }{ 18} + \frac{ 2 }{ 3 } = -\frac{ 2 }{ 18} + \frac{ 12 }{ 18 } = \frac{ -2 + 12 }{ 18 } = \frac{ 10 }{18 } = \frac{ 5 }{ 9 }\]

OpenStudy (ranga):

If any step is not clear feel free to ask.

OpenStudy (anonymous):

makes sense